OpenStudy (anonymous):

Simplify the expression. x sqrt 8x^3+x sqrt 2+ 4sqrt2 write in standerd form #ANWSER

OpenStudy (lgbasallote):

\[ x \sqrt{8x^3} + x\sqrt 2 + 4\sqrt 2 \implies x\sqrt{4 \times x^2 \times 2 \times x} + x\sqrt 2 + 4\sqrt 2\] \[\implies 2x^2\sqrt{2x} + x\sqrt 2 + 4\sqrt 2\] does that help?

OpenStudy (anonymous):

naw